OpenStudy (anonymous):

Circle A has an equation of (x + 8)squared + (y + 3)squared = 16. Where is the center of the circle located? in the first quadrant on the y-axis on the x-axis in the third quadrant

OpenStudy (eyust707):

okay so now they have taken it a step further

OpenStudy (eyust707):

You know to find the center and you know how to find the radius. This time they give you an equation that tells you information about the center and information about the radius.

OpenStudy (eyust707):

The standard equation of a circle is as follows: \[(x-h)^2 +(y-k)^2 = r ^2 \]

OpenStudy (eyust707):

The r represents the radius. The h represents the x-value of the center The k represents the y-value of the center

OpenStudy (eyust707):

The x and y just stay as x and y.

OpenStudy (eyust707):

You understand that so far?

OpenStudy (anonymous):

trying to but continue

OpenStudy (eyust707):

So your equation is: \[(x+8)^2 + (y+3)^2 = 16\]

OpenStudy (eyust707):

Look at the standard equation, what would you have to put in for h to get it to look like yours?

OpenStudy (anonymous):

ok i think you have completely lost me :-(

OpenStudy (eyust707):

\[(x-h)^2 +(y-k)^2 = r ^2\] \[(x+8)^2 + (y+3)^2 = 16\]

OpenStudy (eyust707):

look at the botttom equation

OpenStudy (eyust707):

what number is in the place of "h"

OpenStudy (anonymous):

8

OpenStudy (eyust707):

-8

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

center , -8,-3

OpenStudy (eyust707):

-8 is the x value of the center

OpenStudy (anonymous):

but (-8,-3) are not first quadrant

OpenStudy (anonymous):

oh is -8-3 in the third quadrant?

OpenStudy (eyust707):

yes... sam ruined it...

OpenStudy (eyust707):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

well i have another question that you can help me with

OpenStudy (eyust707):

sure post it

OpenStudy (eyust707):

you can just put it here if you want

OpenStudy (anonymous):

A circle is shown on the coordinate plane below. What is the equation of this circle? (x – 1)2 + y2 = 2 (x + 1)2 + y2 = 4 (x + 1)2 + y2 = 2 (x – 1)2 + y2 = 4

OpenStudy (eyust707):

Whats the radius of the circle shown?

OpenStudy (anonymous):

2?

OpenStudy (eyust707):

Yep! and the center?

OpenStudy (anonymous):

0,1?

OpenStudy (eyust707):

the x value always comes first...

OpenStudy (eyust707):

(1,0)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok

OpenStudy (eyust707):

So: r=2 center= (1,0) That means h= 1 k = 0

OpenStudy (eyust707):

Now just plug r h and k into the standard equation. \[(x-h)^2 +(y-k)^2 = r ^2\]

OpenStudy (anonymous):

lol don't worry sam isn't gonna ruin this 1

OpenStudy (anonymous):

lol sam

OpenStudy (anonymous):

so r is 2 and h is -1?

OpenStudy (eyust707):

be careful when you plug in